# Mixed-State Quantum Denoising Diffusion Probabilistic Model

### v1

@@ I Introduction
Classical deep generative models (DGMs) such as generative adversarial network (GAN) [8], variational autoencoder (VAE) [9], and attention-based models like Transformers [10] have substantially advanced machine learning by producing realistic samples across various domains such as vision [11], language [12], speech [13], and audio generation [14]. Among the plethora of versatile generative models, denoising diffusion probabilistic models (DDPMs) and their variations [15, 16, 17], which represent a branch of hierarchical VAEs incorporating a thermodynamic approach in deep learning [18, 19], have provided another way to create realistic samples. The model includes two main components. The forward process gradually injects Gaussian noise into the data and transforms it into complete Gaussian white noise. The backward process learns from intermediate forward samples to effectively remove noise using deep neural network.
@@ II.2 Backward denoising process
The backward denoising process begins with an ensemble $\{\tilde{\rho}_{T}\}$ of $n_{\text{train}}$ copies of $n$-qubit maximally mixed states and gradually produces $\{\tilde{\rho}_{0}\}$ that resembles the original state ensemble $\{\rho_{0}\}$. This process utilizes a VQA with $T$ total chunks of sequentially trainable VQCs and projective measurements with $n_{a}$ ancillary qubits [25]. We considered two initialization methods for the ancillary qubits, one uses all zero states, ($|0\rangle^{\otimes n_{a}}$), and the other replaces a single qubit with a Haar random pure state, ($|\phi_{\rm Haar}\rangle\otimes(|0\rangle)^{\otimes(n_{a}-1)}$), to introduce randomness and therefore enhance the generative power. We sample the $n_{\text{train}}$ Haar random states for each diffusion step and maintain these ancillary states throughout the individual training process. After training, we freshly draw the states for sampling and input generation in subsequent stages. The final quantum state ensemble $\{\tilde{\rho}_{0}\}$ is produced by running the sequence of trained PQCs in reverse order, using $n_{\rm test}$ samples of maximally mixed states.
Having the forward and backward quantum circuits, Fig. 1(b) illustrates how training proceeds at stage $t=m+1$. The forward depolarizing model generates $n_{\text{train}}$ samples in the target ensemble $\{\rho_{m}\}$ with $\Phi_{1},\Phi_{2},\cdots,\Phi_{m-1}$, and $\Phi_{m}$, while the backward PQC outputs the ensemble $\{\tilde{\rho}_{m}\}$ using $\tilde{U}_{T},\tilde{U}_{T-1},\cdots,\tilde{U}_{m+2}$, and $\tilde{U}_{m+1}$. Based on the loss function that captures the resemblance between the output ensemble $\{\tilde{\rho}_{m}\}$ and the forward depolarizing result $\{\rho_{m}\}$, the gradient-based algorithm trains the parameters $\bm{\theta}_{m+1}$ of the PQC. After optimization, the parameters $\bm{\theta}_{m+1}$ are fixed and reused in further steps. By utilizing this approach, the model can maintain a sufficiently small number of trainable parameters to ensure convergence and avoid training issues known as the barren plateau [30, 31].
@@ III Applications
There are several tasks utilizing MSQuDDPM to generate the desired quantum state ensembles. In this section, we present these tasks, their results, and interpretations. Since the initial data for training and testing of the backward process are all maximally mixed states, we use the forward diffusion and backward testing samples for visualization. Details of the simulations are provided in Table 1.
@@ III.1 Clustered states
In the clustered state generation task, the initial state ensemble is the set of density matrices $\{\rho_{0}\}$, where each $\rho_{0}$ is defined as $\rho_{0}=(1-q_{0})|\psi_{0}\rangle\langle\psi_{0}|+q_{0}I/d$. Here, $q_{0}\in[0,0.01)$ is the depolarizing parameter, $I/d$ represents the maximally mixed state, and $|\psi_{0}\rangle=|0\rangle+\epsilon_{0}c_{0}|1\rangle$ up to normalization coefficient with $\epsilon_{0}=0.08$, and ${\rm Re}(c_{0}),{\rm Im}(c_{0})\sim\mathcal{N}(0,1)$. We employ two ancillary qubits, initialized as $|\phi_{\rm Haar}\rangle\otimes|0\rangle$. The training of the backward PQC uses the Wasserstein distance cost function with superfidelity between the backward circuit outputs at step $t$, $\{\tilde{\rho}_{t-1}\}$, and forward diffusion results at step $t-1$, $\{\rho_{t-1}\}$. Fig. 1(a) and (c) depict samples from the forward and backward processes, illustrating MSQuDDPM’s ability to recover the diffusion process of the clustered distribution.
@@ IV.2 Noise scheduling in MSQuDDPM
In multi-qubit systems, the outputs of the MSQuDDPM forward process with linear or cosine schedules tend to quickly lose their original characteristics, such as purity or magnetization. Fig. 5(a) illustrates the decay of average purity, defined as $\sum_{\rho\in\{\rho_{t}\}}\text{Tr}(\rho^{2})/|\{\rho_{t}\}|$, during the forward depolarizing process under different forward scheduling methods. The purities of linear and cosine schedules rapidly converge to the maximally mixed states by $t=3$. Fast forgetting of the original form exacerbates the MSQuDDPM’s performance, as it makes not only the later training rounds hard to converge but also the earlier training steps ineffective because the target state ensemble gathers close to the maximally mixed states instead of spreading out for better optimization. Thus, maintaining the original state is crucial when handling multi-qubit ensemble generation tasks with MSQuDDPM. To address these constraints, cosine-exponent scheduling, especially cosine square scheduling, is devised to better conserve the initial quantum states. In Fig. 5(a), we can observe that cosine square scheduling produces a smoother transition in average purity compared to other scheduling methods in the many-body phase learning task. In Fig. 5(b), we compare the MMD distance between the intermediate training results $\{\tilde{\rho}_{t}\}$ and the samples from the initial distribution $\{\rho_{0}\}$ across models using different noise scheduling methods in the same task. The results show significant differences in performance across the noise scheduling methods (see Table 1 for numerical details). The model with cosine square scheduling successfully recovers the original ensemble, whereas the other models struggle to converge.

### v2

@@ I Introduction
Classical deep generative models (DGMs) such as generative adversarial network (GAN) [8], variational autoencoder (VAE) [9], and attention-based models, including Transformers [10], have substantially advanced machine learning by producing realistic samples across various domains such as vision [11], language [12], speech [13], and audio generation [14]. Among the plethora of versatile generative models, denoising diffusion probabilistic models (DDPMs) and their variations [15, 16, 17], which represent a branch of hierarchical VAEs incorporating a thermodynamic approach in deep learning [18, 19], have provided another way to create realistic samples. The model includes two main components. The forward process gradually injects Gaussian noise into the data and transforms it into complete Gaussian white noise. The backward process learns from intermediate forward samples to effectively remove noise using deep neural network.
@@ II.2 Backward denoising process
The backward denoising process begins with an ensemble $\{\tilde{\rho}_{T}\}$ of $n_{\text{train}}$ copies of $n$-qubit maximally mixed states and gradually produces $\{\tilde{\rho}_{0}\}$ that resembles the original state ensemble $\{\rho_{0}\}$. This process utilizes a VQA with $T$ total chunks of sequentially trainable VQCs and projective measurements with $n_{a}$ ancillary qubits [25]. We considered two initialization methods for the ancillary qubits, one uses all zero states, ($|0\rangle^{\otimes n_{a}}$), and the other replaces a single qubit with a Haar random pure state, ($|\phi_{\rm Haar}\rangle\otimes(|0\rangle)^{\otimes(n_{a}-1)}$), to introduce randomness and therefore enhance the generative power. We sample the $n_{\text{train}}$ Haar random states for each diffusion step and maintain these ancillary states throughout the individual training process. After training, we freshly draw the states for sampling and input generation in subsequent stages. The final quantum state ensemble $\{\tilde{\rho}_{0}\}$ is produced by running the sequence of trained PQCs in reverse order, using $n_{\rm test}$ samples of maximally mixed states. The circuit layout of the forward and backward process is presented in Appendix A.
Having the forward and backward quantum circuits, Fig. 1(b) illustrates how training proceeds at stage $t=m+1$. The forward depolarizing model generates $n_{\text{train}}$ samples in the target ensemble $\{\rho_{m}\}$ with $\Phi_{1},\Phi_{2},\cdots,\Phi_{m-1}$, and $\Phi_{m}$, while the backward PQC outputs the ensemble $\{\tilde{\rho}_{m}\}$ using $\tilde{U}_{T},\tilde{U}_{T-1},\cdots,\tilde{U}_{m+2}$, and $\tilde{U}_{m+1}$. Based on the loss function that captures the resemblance between the output ensemble $\{\tilde{\rho}_{m}\}$ and the forward depolarizing result $\{\rho_{m}\}$, the gradient-based algorithm trains the parameters $\bm{\theta}_{m+1}$ of the PQC. After optimization, the parameters $\bm{\theta}_{m+1}$ are fixed and reused in further steps. By leveraging this ‘divide-and-conquer’ strategy, MSQuDDPM allows easier optimization at each stage with relatively small discrepancies between consecutive ensembles. Following the QuDDPM’s trainability analysis [25], suppose the model requires at least a polynomial circuit depth in the system size $n$ to generate quantum ensembles. By splitting the problem into multiple subproblems with circuit depth scaling logarithmically with the system size, $\log(n)$, MsQuDDPM can maintain shallow circuits in each step to ensure convergence and avoid training issues such as the barren plateau [30, 31].
@@ III Applications
There are several tasks utilizing MSQuDDPM to generate the desired quantum state ensembles. In this section, we present these tasks, their results, and interpretations. Since the initial data for training and testing of the backward process are all maximally mixed states, we use the forward diffusion and backward testing samples for visualization. Note that the initial parameters of each task are set as examples to evaluate MSQuDDPM’s performance and can be adjusted to accomodate various problem setups. Details of the simulations and implementation are provided in Table 1 and Appendix B, respectively.
@@ III.1 Clustered states
In the clustered state generation task, the initial state ensemble is the set of density matrices $\{\rho_{0}\}$, where each $\rho_{0}$ is defined as $\rho_{0}=(1-q_{0})|\psi_{0}\rangle\langle\psi_{0}|+q_{0}I/d$. Here, $q_{0}\in[0,0.01)$ is the depolarizing parameter to construct the initial ensemble with mixed states, $I/d$ represents the maximally mixed state, and $|\psi_{0}\rangle=|0\rangle+\epsilon_{0}c_{0}|1\rangle$ up to normalization coefficient with $\epsilon_{0}=0.08$, and ${\rm Re}(c_{0}),{\rm Im}(c_{0})\sim\mathcal{N}(0,1)$. We employ two ancillary qubits, initialized as $|\phi_{\rm Haar}\rangle\otimes|0\rangle$. The training of the backward PQC uses the Wasserstein distance cost function with superfidelity between the backward circuit outputs at step $t$, $\{\tilde{\rho}_{t-1}\}$, and forward diffusion results at step $t-1$, $\{\rho_{t-1}\}$. Fig. 1(a) and (c) depict samples from the forward and backward processes, illustrating MSQuDDPM’s ability to recover the diffusion process of the clustered distribution.
@@ IV.2 Noise scheduling in MSQuDDPM
In multi-qubit systems, the outputs of the MSQuDDPM’s forward process with linear or cosine schedules tend to quickly lose their original characteristics, such as purity or magnetization. Fig. 5(a) illustrates the decay of average purity, defined as $\sum_{\rho\in\{\rho_{t}\}}\text{Tr}(\rho^{2})/|\{\rho_{t}\}|$, during the forward depolarizing process under different forward scheduling methods. The purities of linear and cosine schedules rapidly converge to the maximally mixed states by $t=3$. Fast forgetting of the original form exacerbates the MSQuDDPM’s performance, as it makes not only the later training rounds hard to converge but also the earlier training steps ineffective because the target state ensemble gathers close to the maximally mixed states instead of spreading out for better optimization. Thus, maintaining the original state is crucial when handling multi-qubit ensemble generation tasks with MSQuDDPM. To address these constraints, cosine-exponent scheduling, especially cosine square scheduling, is devised to better conserve the initial quantum states. In Fig. 5(a), we can observe that cosine square scheduling produces a smoother transition in average purity compared to other scheduling methods in the many-body phase learning task. A more detailed analysis of purity decay behavior under different noise scheduling methods is presented in Appendix C, where we extend the forward process simulations to other tasks. Our results indicate that while specific characteristics may vary depending on the task, the overall trend of purity decay can be universal across forward schedulings. In Fig. 5(b), we compare the MMD distance between the intermediate training results $\{\tilde{\rho}_{t}\}$ and the samples from the initial distribution $\{\rho_{0}\}$ across models using different noise scheduling methods in the same task. The results show significant differences in performance across the noise scheduling methods (see Table 1 for numerical details). The model with cosine square scheduling successfully recovers the original ensemble, whereas the other models struggle to converge.
@@ V Discussions
Another potential direction is to implement the proposed models on quantum devices and examine the scalability of the approach. The numerical simulations of MSQuDDPM are performed on classical computers, which require exponentially growing space to store and process mixed-state quantum data. Although the simulations demonstrate results up to four qubits due to resource limitations, we anticipate that MSQuDDPM can be applied to larger qubit systems when the individual VQC is designed to be sufficiently shallow, depending on the tasks and input qubits. Thus, deploying the models on quantum devices would provide insights into the method’s scalability.
@@ Appendix B Implementation details
Figure 8: The mean purity decay of samples throughout the forward process of quantum state ensemble generation tasks. We used $n_{\text{train}}=n_{\text{test}}=100$ samples for each task, including (a) clustered states for $n=1$ qubit, (b) circular states for $n=1$ qubit, (c) many-body phase learning for $n=8$ qubits, and (d) many-body phase learning for $n=12$ qubits. In each plot, the gray, green, and purple lines represent the average purity decay under linear, cosine, and cosine square scheduling, respectively.
Appendix C Additional details on noise scheduling in MSQuDDPM
In this section, we analyze the forward process with different noisy scheduling strategies considered in the main text. As shown in Fig. 8, for different target state ensembles, the cosine square scheduling in Eq. (2) enables a nontrivial states purity in more forward steps, which provides learnable information for the backward process. In contrast, the linear and cosine scheduling strategies undergo a sharp decay in the first few steps and make the states to be fully mixed in the late-stage, which in turn restricts the number of effective steps for learning in the backward process and thus results in limited performance as expected shown in Fig. 5. We expect the advantage from forward noise scheduling to be applicable to different learning tasks, and a detailed comparison of learning performance and a strategy for selecting an appropriate value of $k$ in Eq. (2) remain as future work.